OpenStudy (anonymous):

Lolita and Doug working together can paint a shed in 3 hours and 20 minutes. If Doug can paint the shed by himself in 10 hours, how long would it take Lolita to paint the shed by herself?

OpenStudy (anonymous):

If Lolita and Doug can paint the shed in 3 hours and 20 minutes, then that means that they can paint the shed in 10/3 hours. Distance = rate x time: 1 shed = 10/3*r --> r = 3/10 shed/hour. This means they can paint 3/10 of the shed in one hour. If Doug takes 10 hours; then he can paint 1/10 of the shed in an hour. 3/10*t = 1/10*t +Lt, where L is the amount of shed she can paint in one hour. L equals 1/5 if you cancel out the t's. If 1 shed = 1/5*t, then she can paint 1 shed in 5 hours.